One pulley rotates at 560 rpm. A second pulley, connected to the first by a belt, has a diameter of 12 in and rotates at 330 rpm. Find the diameter of the first pulley.
Answer #1

Let , d1 = diameter of driver (first pulley)

n1 = speed of driver

d2 = diameter of follower (second pulley)

n2 = speed of follower

length of the belt passes over the driver in one unit time = \pid1n1

length of the belt passes over the driver in one unit time = \pid2n2

since length of the belt is constant , therefore , \pid1n1 = \pid2n2

​​​​​​   n2/n1 =  d1/d2

   330/560 = d1/12

d1 = (330*12)/560

d1 = 7.1 in

Hence diameter of first pulley is found to be 7.1 in